# Router-Tuning: A Simple and Effective Approach for Dynamic Depth

### v2

@@ Abstract
The Mixture of Depths (MoD) was introduced to dynamically adjust the computational depth by skipping less important layers, thereby reducing computation and improving efficiency. Despite its promise, existing MoD approaches remain under-explored and face two main challenges: (1) high training costs due to the need to train the entire model along with the routers that determine which layers to skip, and (2) the risk of performance degradation when important layers are bypassed. In response to the first issue, we propose Router-Tuning, which fine-tunes only the routers on a small dataset, drastically reducing the computational overhead associated with full model training. For the second challenge, we investigate Router-Tuning across different architectures and granularities, demonstrating its effectiveness on Attention layers and MoE layers. This method preserves the model’s performance while significantly enhancing computational and memory efficiency. Extensive experiments demonstrate that our approach delivers competitive results while dramatically improving the computation efficiency, e.g., 21% speedup and only a 0.2% performance drop. The code will be released at https://github.com/CASE-Lab-UMD/Router-Tuning.
@@ 1 Introduction
Despite its potential, current MoD methods are still underexplored and face several critical challenges. On the one hand, the involvement of additional router networks, which decide which layers to skip, often requires extra extensive training: Raposo et al. (2024) train the entire model from scratch while Tan et al. (2024) performs costly continual training. This creates a significant barrier to efficiently integrating MoD with existing LLMs. Furthermore, most prior MoD implementations Raposo et al. (2024); Tan et al. (2024) have been applied primarily to transformer blocks and MLP layers, which are sensitive to skipping. As a result, omitting important components often leads to significant performance degradation He et al. (2024b).
@@ 3.1 Motivation
The Mixture of Depths (MoD) framework Raposo et al. (2024), which dynamically adjusts layer depth based on input complexity to enhance computational efficiency, was originally designed for integration during the pretraining phase, where transformer models are trained from scratch with MoD-enabled layers. More recently, Tan et al. (2024) applied MoD to pretrained Llama models Touvron et al. (2023) through continual training. While these approaches have shown promise, they are computationally expensive and time-consuming, limiting scalability and practical deployment. A more efficient alternative is to apply MoD directly to existing pretrained models, followed by lightweight fine-tuning of a subset of parameters Houlsby et al. (2019b); Hu et al. (2022), significantly reducing both computational costs and training time.
On the other hand, MoD has typically been implemented at the transformer block level. However, skipping entire transformer blocks has shown to be suboptimal to maintain the performance. Inspired by He et al. (2024b), we recognize that each transformer block contains layers of varying importance. Aggressively skipping entire blocks risks omitting critical layers, potentially degrading performance. Instead, skipping fine-grained layers offers a more effective strategy for preserving model accuracy. Moreover, unlike blocks that generally share the same architecture, individual layers impose different computational costs. For instance, in dense transformer models, attention layers are particularly expensive, with computational complexity scaling quadratically with sequence length and additional memory needed for KV cache storage. In contrast, in MoE models, MLP layers house the majority of the parameters, leading to substantial communication and computation overhead. These observations motivate us to target fine-grained layers for dynamic depth adjustments, aiming to achieve a more effective balance between performance and efficiency.
@@ 3.2 Router-Tuning for Dynamic Depth
As shown in Figure 1, Router-Tuning incorporates an additional router that determines whether to skip the layer. Router-Tuning can be deployed in two levels: (1) token-level, where layers are dynamically skipped for individual tokens, and (2) sequence-level, where layers are dynamically skipped for the entire sequence.
Specifically, given an input $\bm{x}\in\mathbb{R}^{L\times d}$, the router first computes an importance score for the input:
$\bm{R}(\bm{x})=\begin{cases}\text{GATE}(\bm{x}_{i}),&\text{Token-level}\\ \text{GATE}(\frac{1}{L}\sum_{i=1}^{L}\bm{x}_{i}),&\text{Sequence-level}\\ \end{cases},$ (1)
where $\bm{R}$ is a scoring router that assesses the importance score of the input, GATE is the gating function $\bm{y}=\text{Sigmoid}(\bm{W}\bm{x})$. Given the computed importance scores, we further utilize a binarized mask $\bm{M}$ to indicate the remaining tokens or sequences:
@@
where $\tau$ is the threshold that determines whether to skip a layer. The score is set to zero for skipped inputs and one for retained inputs, ensuring stable outputs Tan et al. (2024).
To enable a differentiable and trainable binary decision process, we utilize the straight-through estimator (STE) Bengio et al. (2013), which allows gradients to propagate through the binary selection mechanism. The final output of MoD is then computed as follows:
@@
$\frac{\partial\bm{y}}{\partial\bm{R}}=\frac{\partial\bm{y}}{\partial{\bm{M}}}\frac{\partial{\bm{M}}}{\partial{\bm{R}}}\frac{\partial\bm{R}}{\partial{\bm{W}}}.$ (4)
@@
This dynamic routing mechanism ensures that computation is performed only when necessary—specifically, when the score $\bm{R}(\bm{x})$ meets or exceeds the threshold $\tau$—thereby enhancing the computational efficiency.
@@ 3.3 Extension to Mixture of Experts
Mixture of Experts (MoE) employs sparse activation, dynamically selecting expert networks for each input, which significantly reduces computational resource demands. MoE is primarily used to replace MLP layers in transformer blocks, enhancing representational power and delivering promising performance Jiang et al. (2024); Dai et al. (2024); Muennighoff et al. (2024). However, MoE layers impose substantial memory and computational costs, as they multiply the computation required for MLP layers. Moreover, MoE exhibits a high degree of redundancy, allowing certain experts or layers to be skipped with minimal impact on performance Lu et al. (2024); He et al. (2024a). Building on this, we extend Router-Tuning to MoE layers by implementing dynamic skipping within each expert:
@@
where $\bm{E}_{i}$ denotes the $i$-th expert and $\hat{\bm{E}}_{i}(\bm{x})$ is the corresponding output denotes the corresponding output, bypassing the skipped tokenss. Given a collection of $n$ experts, $\{\bm{E}_{1},\bm{E}_{2},\dots,\bm{E}_{n}\}$, the overall output of the MoE layer is as follows:
@@ 3.4 Training Objectives
Specifically, we optimize two training objectives: task-specific performance and MoD capacity (the proportion of non-skipped inputs). On the one hand, Router-Tuning is designed to maintain the performance of the original model, which we enforce using the loss term $\mathcal{L}_{\text{task}}$ during fine-tuning. On the other hand, while maintaining comparable performance, the model is encouraged to skip more tokens (i.e., reduce MoD capacity) to enhance efficiency. To achieve this, we introduce another loss term $\mathcal{L}_{\text{MoD}}$, which drives the model to reduce MoD capacity to a desired target sparsity level $s$, thereby lowering computational costs and accelerating inference. The overall training objective is as follows:
@@ 4 Experiment Setup
We conduct experiments on Llama Touvron et al. (2023); Grattafiori et al. (2024), Qwen Bai et al. (2023), and Mistral Jiang et al. (2023) due to their competitive performance and widespread adoption. Additionally, we leverage OLMoE Muennighoff et al. (2024) and Deepseek-MoE Dai et al. (2024) as the backbone to deploy Router-Tuning on the Mixture of Experts framework.
@@
For the training dataset, we used Llama-Pro Wu et al. (2024), given it spanning general instruction, math, and code for the SFT process and offering a wealth of instruction data with varying complexity levels. To evaluate model performance, we report normalized zero-shot or few-shot accuracy on the LM-Harness benchmark. The number of shots for each task is detailed in Table 1, which includes multiple tasks: ARC-C Clark et al. (2018), BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021), OBQA Mihaylov et al. (2018), PIQA Bisk et al. (2019), RTE Wang et al. (2019), and WinoGrande ai2 (2019). The evaluation code is based on EleutherAI’s LM Harness framework Gao et al. (2023).
@@
Table 2: Experimental results of Router-Tuning deployed at different granularities. While Router-Tuning is primarily applied to Attention layers, we also evaluate its performance on Block and MLP layers for comparison. The number of skippable layers is constrained to 16, and the overall MoD capacity is 50%.
@@
Hyperparameters
We set $\tau$ as 0.5, which corresponds to the midpoint of the sigmoid function. To ensure that training starts from dense models, we initialize $\bm{W}$ to zero, ensuring that $\bm{R}(\bm{x})\geq\tau$ initially, i.e., training from dense models. To achieve the desired MoD capacity, we perform a grid search over the learning rate from {1e-5, 2e-5, 5e-5, 1e-4, 2e-4} and the scale factor $\lambda$ from {0, 0.1, 0.01, 0.001}, respectively.
To evaluate the performance of the model, we report the results of the following tasks: BoolQ Clark et al. (2019), OBQA Mihaylov et al. (2018), PIQA Bisk et al. (2019), RTE Wang et al. (2019), ARC-C Clark et al. (2018), HellaSwagZellers et al. (2019), MMLU Hendrycks et al. (2021), WinoGrande ai2 (2019) and GSM8K Cobbe et al. (2021). Please refer to Table 1 for detailed information. The evaluation code is based on EleutherAI LM Evaluation Harness Gao et al. (2023).
@@ 5 Main Results
In this section, we evaluate the effectiveness of Router-Tuning on mainstream transformer architectures.
5.1 Competitive Performance of Router-Tuning
Router-Tuning achieves superior performance in Attention layers We first compare deploying Router-Tuning to different modules, e.g., Block, MLP, and Attention. Based on the observation that deeper layers are more redundant than shallow layers Gromov et al. (2024); He et al. (2024b), we focus on deploying Router-Tuning to the deepest layers except the last one, leaving other layers unchanged. In Table 2, we transform half of the blocks/layers into MoD.
While previous studies have primarily explored layer dropping or skipping to Block and MLP layers Bae et al. (2023); Gromov et al. (2024), skipping these modules significantly degrades performance when applied at either token or sequence level. In contrast, applying dynamic depth to Attention layers maintains performance close to that of the original models, e.g., 69.4 v.s. 69.7 in Llama-3-8B. These findings reinforce our motivation to focus on Router-Tuning on Attention, and we utilize Router-Tuning on Attention Layers by default unless stated otherwise.
@@ 5.1 Competitive Performance of Router-Tuning
While Attention Drop He et al. (2024b), which statically drops attention layers, has demonstrated the promising performance of Layer Drop, its static nature lacks flexibility and limits representational power. Therefore, we further investigate the improvements offered by the dynamic mechanism. As shown in Figure 3, we compare Router-Tuning with Attention Drop He et al. (2024b) under the same computation budget — for instance, dropping 4 layers versus deploying MoD to 8 layers with 50% capacity. Under the same skipping ratios, Router-Tuning significantly outperforms Attention Drop He et al. (2024b) on the GSM8K benchmark Cobbe et al. (2021), e.g. 6.5% when the skipping ratio is 25.0%. In Figure 2, we further visualize the layer-wise skipping ratios of MoD versus Attention Drop. Unlike static approaches that permanently remove certain layers, Router-Tuning maintains the utilization of all layers by adaptively distributing skipping ratios across them. This flexible allocation strategy contributes to improved performance.
Figure 2: Skipping Ratios for MoD and Attention Drop.
Table 3: Ablation study of Router-Tuning across different models.
@@ 5.2 Efficiency Improvements
We measure the efficiency in both training and inference, focusing on computational and memory usage.
@@
The training efficiency of our method lies in two perspectives: trainable parameters and training steps. Since the router projects the input from dimension $d$ to 1, the number of trainable parameters is $d\times 1$ per layer, and the total number of trainable parameters is fewer than 0.01% of the whole model. Additionally, Router-Tuning only requires a few steps, which is verified in Section 6. Consequently, Router-Tuning on a single Nvidia A6000 GPU only takes less than 15 minutes, which is over 1000 times faster than DLO Tan et al. (2024) that requires 36 hours on Nvidia RTX A100 GPUs.
@@
We also evaluate the run-time speed improvements achieved with Router-Tuning. The inference speed is measured throughout the entire generation process, from the initial input prompt to the generation of the final token. To ensure that our results accurately reflect the performance gains, we adhere to two key principles: (1) all operations are performed on a single Nvidia RTX A6000 Ada GPU, eliminating any communication overhead from multi-GPU setups; and (2) we increase batch sizes to fully utilize the GPU for each model. As shown in Table 2, different granularities contribute to different levels of speedup. This variation is primarily due to attention layers, where fine-grained token-level MoD introduces differing token lengths within a batch, necessitating padding operations to standardize sequence lengths. Instead, the speedups at the sequence level surpass those at the token level, with Router-Tuning achieving a 21% improvement in inference speed.
@@
The KV cache stores intermediate representations of attention layers, accelerating inference by eliminating redundant computations but incurring substantial memory overhead. Our approach, which selectively skips attention layers, significantly reduces KV cache size—for instance, achieving an 8GB reduction when processing an input sequence of length 2048 with a batch size of 64 on Llama-3-8B. In contrast, DLO operates exclusively on MLP layers and retains the full KV cache, providing no memory savings.

### v4

@@ Abstract
The Mixture of Depths (MoD) was introduced to dynamically adjust the computational depth by skipping less important layers, thereby reducing computation and improving efficiency. Despite its promise, existing MoD approaches remain under-explored and face two main challenges: (1) high training costs due to the need to train the entire model along with the routers that determine which layers to skip, and (2) the risk of performance degradation when important layers are bypassed. In response to the first issue, we propose Router-Tuning, which fine-tunes only the routers on a small dataset, drastically reducing the computational overhead associated with full model training. For the second challenge, we investigate Router-Tuning across different architectures and granularities, demonstrating its effectiveness on Attention layers and MoE layers. This method preserves the model’s performance while significantly enhancing computational and memory efficiency. Extensive experiments demonstrate that our approach delivers competitive results while dramatically improving the computation efficiency, e.g., 21% speedup and only a 0.2% performance drop. The code is released at https://github.com/CASE-Lab-UMD/Router-Tuning.
@@ 1 Introduction
Despite its potential, current MoD methods are still underexplored and face several critical challenges. On the one hand, the involvement of additional router networks, which decide which layers to skip, often requires extra extensive training: Raposo et al. (2024) train the entire model from scratch while Tan et al. (2024) performs costly continual training. This creates a significant barrier to efficiently integrating MoD with existing LLMs. Furthermore, most prior MoD implementations Raposo et al. (2024); Tan et al. (2024) have been applied to transformer blocks and MLP layers, which are sensitive to skipping. As a result, omitting important components often leads to significant performance degradation He et al. (2024b).
@@ 3.1 Motivation
The Mixture of Depths (MoD) framework Raposo et al. (2024), which dynamically adjusts layer depth based on input complexity to enhance computational efficiency, was originally designed for integration during the pretraining phase, where transformer models are trained from scratch with MoD-enabled layers. More recently, Tan et al. (2024) applied MoD to pretrained Llama models Touvron et al. (2023) through continual training. While these approaches have demonstrated promising results, deploying MoD remains computationally expensive and time-consuming, posing challenges for scalability and real-world deployment. A more efficient alternative is to apply MoD directly to existing pretrained models, followed by lightweight fine-tuning of a subset of parameters Houlsby et al. (2019b); Hu et al. (2022), significantly reducing both computational costs and training time.
On the other hand, MoD has typically been implemented at the transformer block level. However, skipping entire transformer blocks has shown to be suboptimal to maintain the performance. Inspired by He et al. (2024b), we recognize that each transformer block contains layers of varying importance. Aggressively skipping entire blocks risks omitting critical layers, potentially degrading performance. Instead, skipping fine-grained layers offers a more effective strategy for preserving model accuracy. Moreover, unlike blocks that generally share the same architecture, individual layers impose different computational costs. For instance, in dense transformer models, attention layers are particularly expensive, with computational complexity scaling quadratically with sequence length and additional memory needed for KV cache storage. In contrast, in MoE models, MLP layers hold the majority of the parameters, leading to substantial communication and computation overhead.
These insights motivate us to deploy MoD in a cost-effective manner (i.e., Router-Tuning) while optimizing the balance between performance and efficiency.
@@ 3.2 Router-Tuning for Dynamic Depth
In this part, we propose Router-Tuning to address the challenges outlined in Section 3.1. As illustrated in Figure 1, Router-Tuning incorporates an additional trainable router that determines whether to skip the layer. Specifically, Router-Tuning can be deployed in two levels: (1) token-level, where layers are dynamically skipped for individual tokens, and (2) sequence-level, where layers are dynamically skipped for the entire sequence. Given an input $\bm{x}\in\mathbb{R}^{L\times d}$, the router first computes an importance score for the input:
$\bm{R}(\bm{x})_{i}=\begin{cases}\text{GATE}(\bm{x}_{i}),&\text{Token-level}\\ \text{GATE}(\frac{1}{L}\sum_{i=1}^{L}\bm{x}_{i}),&\text{Sequence-level}\\ \end{cases},$ (1)
where $\bm{R}$ is a scoring router that assesses the importance score of the input, GATE is the gating function $\text{GATE}(\bm{x})=\text{Sigmoid}(\bm{W}\bm{x})$. Based on the computed importance scores, we further apply a binarized mask $\bm{M}$ to determine whether to skip a token or an entire sequence:
@@
where $\tau$ is the threshold. The score is set to zero for skipped inputs and one for retained inputs, ensuring stable outputs Tan et al. (2024).
To enable a differentiable and trainable binary decision process, we utilize the straight-through estimator (STE) Bengio et al. (2013), which allows gradients to propagate through the binary selection mechanism via $\frac{\partial{\bm{M}}}{\partial{\bm{R}}}=1$. The final output of MoD is then computed as follows:
@@
$\frac{\partial\bm{y}}{\partial\bm{W}}=\frac{\partial\bm{y}}{\partial{\bm{M}}}\frac{\partial{\bm{M}}}{\partial{\bm{R}}}\frac{\partial\bm{R}}{\partial{\bm{W}}}.$ (4)
@@
This dynamic routing mechanism ensures that computation is performed only when necessary, thereby enhancing the computational efficiency.
@@ 3.3 Extension to Mixture of Experts
Mixture of Experts (MoE) employs sparse activation, dynamically selecting expert networks for each input, which delivers promising performance in various tasks Jiang et al. (2024); Dai et al. (2024); Muennighoff et al. (2024). However, MoE also exhibits significant redundancy, allowing certain experts or layers to be skipped with minimal impact on performance Lu et al. (2024); He et al. (2024a). Building on this, we extend Router-Tuning to MoE layers by implementing dynamic skipping within each expert:
@@
where $\bm{E}_{i}$ denotes the $i$-th expert and $\hat{\bm{E}}_{i}(\bm{x})$ is the corresponding output denotes the corresponding output, bypassing the skipped tokens. Given a collection of $n$ experts, $\{\bm{E}_{1},\bm{E}_{2},\dots,\bm{E}_{n}\}$, the overall output of the MoE layer is as follows:
@@ 3.4 Training Objectives
Specifically, we optimize two training objectives: improving task-specific performance and lowering MoD capacity (the proportion of non-skipped inputs). On the one hand, Router-Tuning is designed to maintain the performance of the original model, which we enforce using the loss term $\mathcal{L}_{\text{task}}$ during fine-tuning. On the other hand, the model is encouraged to skip more tokens or sequences (i.e., reduce MoD capacity) to enhance efficiency. To achieve this, we introduce another loss term $\mathcal{L}_{\text{MoD}}$, which drives the model to reduce MoD capacity to a desired target sparsity level $s$, thereby lowering computational costs and accelerating inference. The overall training objective is as follows:
@@ 4 Experiment Setup
We conduct experiments on Llama Touvron et al. (2023); Grattafiori et al. (2024), Qwen Bai et al. (2023), and Mistral Jiang et al. (2023) due to their competitive performance and widespread adoption. Additionally, we leverage OLMoE Muennighoff et al. (2024) and Deepseek-MoE Dai et al. (2024) as the backbone to deploy Router-Tuning on the Mixture of Experts.
@@
For the training dataset, we used Llama-Pro Wu et al. (2024), given it spanning general instruction, math, and code for the SFT process and offering a wealth of instruction data with varying complexity levels. To evaluate model performance, we report normalized zero-shot or few-shot accuracy on the LM-Harness benchmark. The number of shots for each task is detailed in Table 1, which includes multiple tasks: ARC-C Clark et al. (2018), BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021), OBQA Mihaylov et al. (2018), PIQA Bisk et al. (2019), RTE Wang et al. (2019), WinoGrande ai2 (2019) and GSM8K Cobbe et al. (2021). The evaluation code is based on EleutherAI’s LM Harness framework Gao et al. (2023).
@@
Hyperparameters
We set $\tau$ as 0.5, which corresponds to the midpoint of the sigmoid function. To ensure that training starts from dense models, we initialize $\bm{W}$ to zero, ensuring that $\bm{R}(\bm{x})\geq\tau$ initially, i.e., training from dense models. To achieve the desired MoD capacity, we perform a grid search over the learning rate from {1e-5, 2e-5, 5e-5, 1e-4, 2e-4} and the scale factor $\lambda$ from {0, 0.1, 0.01, 0.001}, respectively
Table 2: Experimental results of Router-Tuning deployed at different granularities, assessing both task performance and inference efficiency. While Router-Tuning is primarily applied to Attention layers, we also evaluate its performance on Block and MLP layers for comparison. The number of skippable layers is constrained to 16, and the overall MoD capacity is 50%.
@@ 5 Main Results
In this section, we evaluate the effectiveness of Router-Tuning on transformer architectures with the deployment details in Appendix 4.
5.1 Performance of Router-Tuning
Router-Tuning achieves superior performance in Attention layers We first compare deploying Router-Tuning to different modules, e.g., Block, MLP, and Attention, as shown in Table 2. Based on the observation that deeper layers are more redundant than shallow layers Gromov et al. (2024); He et al. (2024b), we focus on deploying Router-Tuning to the deepest layers except the last one, leaving other layers unchanged.
While previous studies have primarily explored layer dropping or skipping to Block and MLP layers Bae et al. (2023); Gromov et al. (2024), skipping these modules significantly degrades performance when applied at either token or sequence level. In contrast, applying dynamic depth to Attention layers maintains the performance of original models, e.g., 69.4 v.s. 69.7 in Llama-3-8B. These findings reinforce our motivation to target Attention layers, and we utilize Router-Tuning on Attention layers by default unless stated otherwise.
Figure 3: Layer-wise Skipping Ratios for MoD and Attention Drop under overall skipping ratios of $50\%$.
@@ 5.1 Performance of Router-Tuning
While Attention Drop He et al. (2024b), which statically drops attention layers, has demonstrated the promising performance of Layer Drop, its static nature lacks flexibility and limits representational power. Here, we further investigate the improvements offered by the dynamic mechanism. As shown in Figure 2, we compare Router-Tuning with Attention Drop He et al. (2024b) under the same computation budget — for instance, dropping 4 layers versus deploying MoD to 8 layers with 50% capacity. Under the same skipping ratios, Router-Tuning significantly outperforms Attention Drop He et al. (2024b) on the GSM8K benchmark Cobbe et al. (2021), e.g. 6.5% when the skipping ratio is 25.0%. In Figure 3, we further visualize the layer-wise skipping ratios of MoD versus Attention Drop. Unlike static approaches that permanently remove certain layers, Router-Tuning maintains the utilization of all layers by adaptively distributing skipping ratios across them. This flexible allocation strategy contributes to improved performance.
Table 3: Ablation study of Router-Tuning across different models architectures, including Llama Touvron et al. (2023), Qwen Bai et al. (2023), Mistral Jiang et al. (2023).
@@ 5.2 Efficiency Improvements
In this part, we measure the efficiency in both training and inference, focusing on computational and memory usage.
@@
The training efficiency of our method lies in two perspectives: trainable parameters and training steps. Since the router projects the input from dimension $d$ to 1, the number of trainable parameters is $d\times 1$ per layer, and the total number of trainable parameters is fewer than 0.01% of the whole model. Additionally, Router-Tuning only requires a few steps, which is verified in Section 6. Consequently, Router-Tuning on a single Nvidia A6000 GPU only takes less than 15 minutes, which is over 1000 times faster than DLO Tan et al. (2024) that finetunes the entire model and requires 36 hours on Nvidia RTX A100 GPUs.
@@
We also evaluate the run-time speed improvements achieved with Router-Tuning. The inference speed is measured throughout the entire generation process, including prefilling and generation. To ensure that our results accurately reflect the performance gains, we adhere to two key principles: (1) all operations are performed on a single Nvidia RTX A6000 Ada GPU, eliminating any communication overhead from multi-GPU setups; and (2) we increase batch sizes to fully utilize the GPU for each model. As shown in Table 2, different granularities contribute to different levels of speedup. This variation is primarily due to attention layers, where fine-grained token-level MoD introduces differing token lengths within a batch, necessitating padding operations to standardize sequence lengths. Instead, the speedups at the sequence level surpass those at the token level, with Router-Tuning achieving a 21% improvement in inference speed.
@@
The KV cache stores intermediate representations of attention layers, accelerating inference by eliminating redundant computations but incurring substantial memory overhead. Our approach, which selectively skips attention layers, significantly reduces KV cache size—for instance, achieving an 8GB reduction when processing an input sequence of length 2048 with a batch size of 64 on Llama-3-8B. In contrast, DLO Tan et al. (2024) operates exclusively on MLP layers and retains the full KV cache, providing no memory savings.
Figure 4: Ablation study on the number of MoD layers.